A 750kg car is travelling along a straight horizontal road at 48km/hr. Calculate the magnitude of constant braking force that will stop the car at a distance 30m.
1
Expert's answer
2020-07-06T15:40:54-0400

The change of kinetic energy is equal to work done by the braking force

"\frac{mv^2}{2}=Fd"

Hence, the braking force

"F=\frac{mv^2}{2d}=\frac{750\times (48\div3.6)^2}{2\times 30}=2222\:\rm N"